OpenStudy (anonymous):

In a certain candy store, 3 pounds of candy and 2 pounds of mints cost $10.80, and 1 pound of candy and 3 pounds of mints cost $5.35. What is the cost per pound of the mints? mints are 65¢ per lb mints are 75¢ per lb mints are 85¢ per lb

OpenStudy (anonymous):

how to solve

OpenStudy (ivettef365):

first make the 2 equations as follows: x = candy y = mints

OpenStudy (ivettef365):

then you have 3x + 2y = 10.80 x + 3y = 5.35

OpenStudy (ivettef365):

can you solve for y ???

OpenStudy (anonymous):

ill try

OpenStudy (ivettef365):

ok let me know if I can help you

OpenStudy (anonymous):

k

OpenStudy (anonymous):

acacia can u help

OpenStudy (ivettef365):

ok what you need to do is multiply (-3) to the second equation so you have 3x +2y = 10.80 x + 3y = 5.35 (-3) then what do you get ?

OpenStudy (anonymous):

found it thx ivettef365

OpenStudy (ivettef365):

what did you get ??